Convert the vector A = xy ax - 2x ay to cylindrical system.


1
Expert's answer
2021-02-11T17:11:03-0500

In cylindrical coordinate system


"x=\rho \cos\phi \\y=\rho \sin\phi \\z=z"

And also


"a_x\n\u200b\t\n =\\cos\u03d5a _\n\u03c1\n\u200b\t\n \u2212\\sin\u03d5a _\n\u03d5\n\u200b\t\n \n\n\\\\a_y=\\sin\\phi a_\\rho +\\cos\\phi a_\\phi \\\\"

Putting all value we get


"A=(\u03c1\\cos\u03d5\u22122)\u03c1 ^\n2\n \\sin\u03d5\\cos\u03d5a _\n\u03c1\n\u200b\t\n\\\\ \u2212(\u03c1\\sin ^\n2\n \u03d5+2\\cos\u03d5)\u03c1\\cos\u03d5a _\n\u03d5\n\u200b"